OpenStudy (anonymous):

Jerry has just received an inheritance of $25,000, and he would like to use it to help him with his retirement. Since Jerry is 25 years old, he figures that the $25,000 can be invested for 40 years before he will need to use it for retirement. Jerry wants to know what interest rate would be necessary for the $25,000 to grow to provide an amount so that he can have a monthly income of $7,000 earned from the interest only. Assume that Jerry is able to invest in real estate and can earn a 15% return on his investment

OpenStudy (dumbcow):

i am getting an interesting equation: the accumulation over 40 years looks like this \[B = 25,000(1+r)^{40}\] the balance needed to obtain monthly interest of 7000 is: \[B(\frac{r}{12}) = 7,000 \rightarrow B=\frac{84,000}{r}\]

OpenStudy (dumbcow):

http://www.wolframalpha.com/input/?i=r%281%2Br%29^40+%3D+84%2F25 so looks like neccessary interest rate is about 9.36%

OpenStudy (anonymous):

sorry,it didnt help

OpenStudy (dumbcow):

what do you mean, thats the solution....9.36% \[25,000(1.0936)^{40} \approx896,000\] monthly interest from that total is about 7,000

OpenStudy (anonymous):

as u r saying 25000*.0936*4=9360

OpenStudy (dumbcow):

?? im sorry i don;t know why you are multiplying by 4

OpenStudy (anonymous):

sorry it was 40, the number of years of deposit. to get the amount after 40 years we will multiply 25000 to 40(no. of years) and to the intersert rate. the interest rate u r telling does not give final amount enough to get a monthly income of $7000.

OpenStudy (dumbcow):

no thats wrong..if you leave the 25,000 in the bank for 40 years then each year it accumulates interest by a factor of (1+r) where r is int rate

OpenStudy (anonymous):

u r absolutely right bt can u show me how the int rate of 9.36 o the $25000 for 40 years gives us $7000 per month

OpenStudy (dumbcow):

what do you mean, thats the solution....9.36% 25,000(1.0936)^40≈896,000 \[896,000*(.0936/12) \approx7,000\]

OpenStudy (anonymous):

i am messed up completely. bt thanks a lot for ur help and time. where are u from ?

OpenStudy (dumbcow):

your welcome and from southwest USA

OpenStudy (anonymous):

i am from Pakistan. i am nt so dumped bt i am messing up with ths que so much that i cant find the right method to solve it

OpenStudy (anonymous):

r u in college or uni?